OpenStudy (anonymous):

Solve "cos (2x) = sin (x) - 2" for [0,2Pi)

sam (.sam.):

cos (2x) = sin (x) - 2 Convert to 1-2sin^2(x) 1-2sin^2(x)= sin (x) - 2 2sin^2(x)+sin(x)-3=0 (sin(x)-1)(2sin(x)+3)=0 sin(x)=1 sin(x)=-3/2 Then use this to solve for x 

OpenStudy (anonymous):

How would you use that chart? I always have trouble negotiating positive/negative values in trig, and it's a very important concept in advanced trig.

OpenStudy (anonymous):

According to the graph, the solution is pi/2

OpenStudy (anonymous):

But I would like to know how this works algebraically

sam (.sam.):

 sin(x)=1 sin(x)=-3/2 Doing sin(x)=1, \[x=\sin^{-1}(1)\] x=90 Then you need to get another value from the chart This sin(x)=1 is positive, so you look at the positive side so, 180-90 and get another value x=90 but its the same. ---------------------------------------- We move on to sin(x)=-3/2 \[x=\sin^{-1}(-3/2)\] No solution . ----------------------------------------- x=90 only